# How do Large Language Models Handle Multilingualism?

### v1

@@ 4.1 Experimental Setup
To comprehensively understand how LLMs work with different abilities, we employ four different kinds of tasks including MGSM (Shi et al., 2022) for reasoning task, XQuAD (Artetxe et al., 2020) for NLU task, XLSum (Hasan et al., 2021) for NLG task, and X-CSQA (Lin et al., 2021) for knowledge question answering task. For XLSum, we randomly sample $500$ data points from the whole test set for each language, while for other tasks, we employ the whole test set.
Original Results
We first evaluate the vanilla performance of Vicuna and Mistral on these datasets for later comparison as presented in Table 3.
Task En De Fr Zh Es Ru
Vicuna XQuAD $57.5$ $50.3$- $55.7$ $55.7$-
MGSM $20.4$ $14.8$ $14.8$ $12.8$ $13.2$ $10.0$
X-CSQA $57.8$ $43.8$ $40.1$ $43.2$ $44.3$ $26.0$
XLSum $13.1$- $14.2$ $61.1$ $10.4$ $20.8$
Mistral XQuAD $57.1$ $48.5$- $64.3$ $54.1$-
MGSM $46.0$ $21.2$ $26.0$ $31.6$ $31.2$ $21.6$
X-CSQA $61.7$ $40.0$ $40.4$ $47.1$ $45.7$ $14.1$
XLSum $13.5$- $15.2$ $56.4$ $10.6$ $21.0$
Table 3: Assessing the baseline performance of Vicuna and Mistral across four representative multilingual tasks in selected languages.
Experiment Details
For reasoning, NLU, and knowledge question answering tasks, we adopt accuracy as the metric. As for the NLG tasks, we adopt ROUGE-L as the metric. Furthermore, for the concrete numbers of different layers, we tune hyperparameters by XQuAD in Chinese. Details are explained in Appendix B.
@@
Tables 4 to 7 present the results of deactivating certain neurons, where “Under” denotes the understanding layers, “S-ATTN” and “S-FFN” correspond to the self-attention and the feed-forward within the task-solving layers respectively, “Gen” refers to the generation layers. The term “Random” is used to describe deactivating randomly chosen neurons, whereas “Lang-Spec” refers to the deactivation of language-specific neurons. We also present the gap between the original performance and performance after deactivation for English ($\Delta_{\text{Eng}}$) and averaged non-English languages ($\Delta_{\text{n-Eng}}$), respectively. A single metric $\Delta$ is then introduced as $\Delta_{\text{Eng}}-\Delta_{\text{n-Eng}}$, where a high value indicates such deactivation operation does not bring much impact to the English performance but lead to performance drop in non-English.
4.2 Analysis on Understanding
@@
Table 4: Results of the understanding task, where ‘✓’ indicates that neurons in the corresponding layer are deactivated, and ‘✗’ signifies they are not. $\Delta$ is defined as the difference between the reduction in performance in English, denoted as $\Delta_{\text{Eng}}$, and the reduction in performance in non-English languages, denoted as $\Delta_{\text{n-Eng}}$.
@@
Table 5: Results of the reasoning task. The highest performance reduction difference ($\Delta$) is achieved by disabling all language-specific neurons, except for those involved in self-attention structure within the task-solving layer.
@@
Table 6: Results of the knowledge question answering. The highest performance reduction difference ($\Delta$) is achieved by disabling all language-specific neurons in the feed-forward structure within the task-solving layer.
@@
Table 7: Results of the generation task following neuron deactivation. The highest performance reduction difference ($\Delta$) is achieved by disabling all language-specific neurons in the generation layer.
Deactivating Method
Table 4 shows the results of the understanding task following the deactivation of five distinct sets of neurons: (i) neurons randomly selected from the understanding layers; (ii) neurons randomly chosen across all layers; (iii) language-specific neurons within the understanding layers; (iv) language-specific neurons in the task-solving layers; (v) language-specific neurons in the generation layers. For a fair comparison, we ensure the numbers of deactivated neurons in all settings are the same. Full results of each language are listed in Appendix C.
@@
We find that by deactivating randomly sampled neurons, no matter in the understanding layer or all layers, the performance of LLMs in both English and non-English languages is almost unaffected compared to other deactivating methods. Note that in some cases, deactivating randomly sampled neurons may even increase the performance because irrelevant neurons are removed, which also aligns with the finding from Sharma et al. (2023). When assessing the differential impact on English and non-English language performance post-deactivation, specifically the difference calculated as $\Delta_{\text{Eng}}-\Delta_{\text{n-Eng}}$, it is evident that the deactivation of random neurons within the understanding layer amplifies this effect. This observation lends partial support to the hypothesized role of the understanding layer in language processing.
Furthermore, we find that deactivating language-specific neurons in the understanding layer influences the performance in English a little while significantly decreasing the performance in non-English languages. When deactivating language-specific neurons in the task-solving layer, both English and non-English languages are significantly reduced while deactivating language-specific neurons in the generation layer influences a little for both English and non-English languages. Therefore, we prove that the first several layers are responsible for understanding because deactivated neurons just disable LLMs on the NLU task in non-English languages. Furthermore, disabling language-specific neurons in the task-solving layer shows that LLMs rely on English, as their performance drops across all languages.
4.3 Analysis on Reasoning
Deactivating Method
Table 5 shows the result of the reasoning task, where we deactivate $6$ sets of neurons. Details are listed in Appendix C.
Findings
4.4 Analysis on Knowledge Question Answering
Deactivating Method
Table 6 shows the result of the knowledge question answering task, where we deactivate $5$ sets of neurons. Details are listed in Appendix C.
Findings
Likewise, the deactivation of randomly selected neurons has a lesser impact compared to language-specific neurons, validating the efficiency of PLND in identifying neurons pertinent to a particular language. Targeted deactivation of language-specific neurons within the feed-forward structure of the task-solving layer predominantly affects non-English languages. This implies that processing multilingual queries necessitates accessing the multilingual information embedded within the relevant structures. However, disabling the self-attention structure compromises the ability to solve tasks across all languages.
4.5 Analysis on Generation
Deactivating Method
Table 7 shows the result of the generation task, where we deactivate $3$ sets of neurons. Details are listed in Appendix C.
Findings
@@ 4.5 Analysis on Generation
Similar to other tasks, the disabling of language-specific neurons within the generation layer attenuates their capacity to generate content in the respective languages. By selectively deactivating neurons that are not associated with English, we do not completely eliminate the models’ multilingual generation abilities. However, as demonstrated in Table 2, the complete deactivation of all language-specific neurons results in the total loss of the LLMs’ multilingual generation capabilities.
5 Further Analysis
5.1 Enhance Multilingual Ability
@@
We have verified the proposed framework for explaining the multilingual working mechanism of LLMs in the above section via deactivating certain neurons. While opposite to employing deactivation, we can also enhance their multilingual ability, especially the understanding and generating ability, by fine-tuning these language-specific neurons.
We employ Llama2-7b-base model for enhancement to eliminate the interference of instruction fine-tuning. We select causal language modeling as our fine-tuning task and create a dataset comprising $200$ documents for each language, extracted from the Wikipedia corpus. [^4] https://dumps.wikimedia.org/ It is important to note that our enhancements are focused on augmenting the model’s capabilities in understanding and generation only; we do not extend its reasoning faculties or broaden its knowledge base as it may require more specific data preparation. Accordingly, we evaluate the efficacy of our enhancements through targeted understanding and generation tasks. Detailed experiment results are shown in Table 8.
Our findings indicate that with just $10$ minutes of fine-tuning on $200$ contexts, LLMs exhibit significant enhancements in multilingual understanding and generation abilities. Notably, there is a relative performance boost of $7.4\%$ on the XQuAD benchmark. Similarly, for XLSum, we observe an $8.9\%$ relative improvement in performance.
Task Method De Fr Zh Es Ru
XQuAD Original $42.2$- $47.3$ $39.6$-
Enhance $\mathbf{45.3}$- $\mathbf{49.7}$ $\mathbf{43.7}$-
XLSum Original -- $8.5$ $48.2$ $9.1$ $19.7$
Enhance -- $\mathbf{9.4}$ $\mathbf{51.6}$ $\mathbf{11.0}$ $\mathbf{21.1}$
Table 8: Enhancement is achieved by fine-tuning the Llama2-7b-base model through a causal language modeling task, utilizing $200$ contexts from each language.
5.2 Analysis on Different Multilingual LLMs
To further verify our proposed framework, we examine two more types of multilingual LLMs, including BLOOMZ (Muennighoff et al., 2022), a hyper-multilingual LLM claiming to support 46 languages, and Chinese Llama Cui et al. (2023), a bilingual LLM focusing on English and Chinese.
Hyper-Multilingual LLMs
Figure 4: Overlapping ratio of language-specific neurons in BLOOMZ.
Bilingual LLMs
We employ Chinese Llama (Cui et al., 2023), which extends existing vocabulary and incorporate secondary pre-training using Chinese data and fine-tune the model with Chinese instruction datasets. However, this intensive training can lead to a degradation in performance for languages other than Chinese. As depicted in Figure 5, Chinese predominates as the primary language for thinking processing and knowledge extraction across all languages. Consequently, the absence of language-specific neurons results in the transformation of it into a Chinese-centric LLM.
Figure 5: Distribution of languages among layers in Chinese Llama given non-English instructions.
@@ 6 Related Work
Interpretability
Conventional interpretability research investigates the significance of input features with their corresponding outputs (Vig, 2019; Hewitt and Liang, 2019; Qiu et al., 2020).In the era of LLMs, one brunch of work includes efforts to understand knowledge storage, with Geva et al. (2021) initiating the study of the feed-forward layer as a knowledge base. Subsequent work has furthered this by altering neuron values (Dai et al., 2022b), mapping embeddings to words (Geva et al., 2022), modifying inputs to recover embeddings (Meng et al., 2022), and analyzing attention heads (Li et al., 2023a). Another line of research centers on the self-attention layer, examining its connection to reasoning (Hou et al., 2023; Stolfo et al., 2023; Friedman et al., 2023) by contrasting the reasoning tree based on attention weights.
Multilingualism
Various studies have been undertaken to construct benchmarks (Zhang et al., 2023), enhance performance through translation (Liang et al., 2023; Huang et al., 2023), aligning representations (Nguyen et al., 2023; Salesky et al., 2023), prompting (Li et al., 2023b; Tanwar et al., 2023). These efforts underscore the importance and complexity of enabling LLMs to operate effectively across multiple languages.
@@ 7 Conclusion
In this work, we introduce a hypothesis suggesting that LLMs address multilingualism by first translating queries into English, processing them using English with the help of multilingual knowledge, and then translating the responses back into the original language. To validate this framework, we introduce a novel method for detecting language-specific neurons and conduct extensive ablation studies. These studies involve selectively deactivating various neuron sets to observe their impact on the multilingual capabilities of LLMs. Additionally, we refine the multilingual performance of LLMs by fine-tuning these language-specific neurons, which only take up a small part of the neurons.
@@ Appendix A English and Non-English Tokens
Appendix B Hyper-parameters

### v2

@@ 3.2 Verification Experiment Setup
To comprehensively understand how LLMs work with different abilities, we employ four kinds of tasks including MGSM (Shi et al., 2022) for reasoning task, XQuAD (Artetxe et al., 2020) for understanding task, X-CSQA (Lin et al., 2021) for knowledge question answering task, and XLSum (Hasan et al., 2021) for generation task. Detailed information regarding these datasets and the testing prompts can be found in Appendix D. We adopt $6$ languages including English (En), German (De), French (Fr), Chinese (Zh), Spanish (Es), and Russian (Ru), as their initial performance on those languages is already quite reasonable for observing the multilingual processing mechanism. For XLSum, we randomly sample $500$ data points from the whole test set for each language taking into consideration its long inference time, while for other tasks, we employ the entire test set. We evaluate the vanilla performance of Vicuna and Mistral on these datasets for later comparison as presented in Appendix E. For reasoning, understanding, and knowledge question answering tasks, we adopt accuracy as the metric. As for the generation tasks, we adopt ROUGE-L as the metric.
Deactivation Strategy
We primarily consider two aspects when selecting the deactivation settings: (1) language-specific neurons versus randomly chosen neurons, and (2) the position of neurons, which encompasses four structures. More detailed settings are explained from Section 3.3 to Section 3.6. For the concrete numbers of different layers, we tune hyperparameters by XQuAD in Chinese. Details are explained in Appendix F.
@@
Tables 2 to 5 present the results of deactivating certain neurons, where “Under” denotes the understanding layers, “S-ATTN” and “S-FFN” correspond to the self-attention and the feed-forward structures within the task-solving layers respectively, “Gen” refers to the generation layers. The term “Random” is used to describe deactivating randomly chosen neurons, whereas “Lang-Spec” refers to the deactivation of language-specific neurons. We also present the gap between the original performance (as shown in Table 9) and performance after deactivation (as shown in Table 12 to Table 15) for English ($\Delta_{\text{Eng}}$) and averaged non-English languages ($\Delta_{\text{n-Eng}}$), respectively. A single metric $\Delta$ is then introduced as $\Delta_{\text{Eng}}-\Delta_{\text{n-Eng}}$, where a high value indicates such deactivation operation does not bring much impact to the English performance but lead to performance drop in non-English. Therefore, this provides evidence that the deactivated neurons are language-specific and hold a significant responsibility in executing the corresponding task.
3.3 Verify the Understanding Stage in MWork
Deactivating Method
Table 2 shows the results of the understanding task following the deactivation of five distinct sets of neurons: (i) neurons randomly selected from the understanding layers; (ii) neurons randomly chosen across all layers; (iii) language-specific neurons within the task-solving layers; (iv) language-specific neurons in the generation layers; (v) language-specific neurons in the understanding layers. For a fair comparison, we ensure the numbers of deactivated neurons in all settings are the same. As mentioned above, in order to verify the functionality of the understanding layer (setting v), we compare it with deactivating other types of layers, specifically setting iii for the task-solving layer and setting iv for the generation layer. Full results are listed in Appendix G.
Table 2: Results of the understanding task, where ‘✗’ indicates that chosen neurons in the corresponding layer are deactivated, and ‘✓’ signifies they are activated. $\Delta$ is defined as the difference between the reduction in performance in English, denoted as $\Delta_{\text{Eng}}$, and the reduction in performance in non-English languages, denoted as $\Delta_{\text{n-Eng}}$.
@@ 3.3 Verify the Understanding Stage in MWork
Lang-Spec $56.2$ $48.3$ $-0.9$ $-7.4$ $+6.5$
Findings
We find that by deactivating randomly sampled neurons, no matter in the understanding layer or all layers, the performance of LLMs in both English and non-English languages is almost unaffected compared to other deactivating methods. Note that in some cases, deactivating randomly sampled neurons may even increase the performance because irrelevant neurons are removed, which also aligns with the finding from (Sharma et al., 2023). When assessing the differential impact on English and non-English language performance after the deactivation, specifically the difference calculated as $\Delta_{\text{Eng}}-\Delta_{\text{n-Eng}}$, it is evident that the deactivation of random neurons within the understanding layer amplifies this effect. This observation lends partial support to the hypothesized role of the understanding layer in language processing.
Furthermore, we find that deactivating language-specific neurons in the understanding layer influences the performance in English a little while significantly decreasing the performance in non-English languages. When deactivating language-specific neurons in the task-solving layer, both English and non-English languages are significantly reduced while deactivating language-specific neurons in the generation layer influences a little for both English and non-English languages. Therefore, we prove that the first several layers are responsible for understanding because deactivated neurons just disable LLMs on the NLU task in non-English languages. Furthermore, disabling language-specific neurons in the task-solving layer shows that LLMs rely on English, as performance drops across all languages.
3.4 Verify the Reasoning Structure in MWork
Deactivating Method
Table 3 shows the result of the reasoning task, where we deactivate $6$ sets of neurons. We adhere to the previous logic of selecting deactivation settings, with the exception that we do not conduct an independent experiment on deactivating neurons in the understanding layer, as its functionality has already been verified. Details are listed in Appendix G.
Table 3: Results of the reasoning task. Disabling all language-specific neurons, except for those involved in self-attention structure within the task-solving layer, greatly reduces performance.
@@ 3.4 Verify the Reasoning Structure in MWork
Findings
3.5 Verify the Knowledge Extraction Structure in MWork
Deactivating Method
Table 4 shows the result of the knowledge question answering task, where we deactivate $5$ sets of neurons. Similarly, we exclude the deactivation of neurons in layers that have already been verified and instead concentrate on the self-attention structure and feed-forward structure in the task-solving layer. Details are listed in Appendix G.
Table 4: Results of the knowledge question answering task. The highest performance reduction difference ($\Delta$) is achieved by disabling all language-specific neurons in the feed-forward structure within the task-solving layer.
@@ 3.5 Verify the Knowledge Extraction Structure in MWork
Findings
Likewise, targeted deactivation of language-specific neurons within the feed-forward structure of the task-solving layer predominantly affects non-English languages. This implies that processing multilingual queries necessitates accessing the multilingual information embedded within the relevant structures. However, disabling the self-attention structure compromises the ability to solve tasks across all languages.
3.6 Verify the Generation Structure in MWork
Deactivating Method
Table 5 shows the result of the generation task, where we deactivate $3$ sets of neurons. Since all previous layers have been verified, we solely deactivate neurons in the generation layer and compare them with randomly selected neurons. Details are listed in Appendix G.
Table 5: Results of the generation task. The highest performance reduction difference ($\Delta$) is achieved by disabling all language-specific neurons in the generation layer.
@@ 3.6 Verify the Generation Structure in MWork
Similar to other tasks, the disabling of language-specific neurons within the generation layer diminishes their capacity to generate content in the respective languages. By selectively deactivating neurons that are not associated with English, we do not completely eliminate the models’ multilingual generation abilities. However, as demonstrated in Table 1, the complete deactivation of all language-specific neurons results in the total loss of the LLMs’ multilingual generation capabilities.
4 Multilingual Enhancement with MWork
@@
We have verified MWork for explaining the multilingual working mechanism of LLMs in the above section via deactivating certain neurons. While opposite to employing deactivation, we can also enhance their multilingual ability, especially the understanding and generating ability, by fine-tuning these language-specific neurons. With language-specific neurons comprising only around $0.1\%$ of all parameters, the need for training documents to improve multilingual capabilities can be significantly reduced to just a few hundred. Additionally, fine-tuning only the language-specific neurons for a particular language does not impact performance in other languages, allowing us to enhance specific languages while preserving performance in others.
MWork helps with enhancing multilingual ability by hundreds of documents.
We employ Mistral-7b-v0.1 for enhancement to eliminate the interference of instruction fine-tuning, and select causal language modeling as our training task. We create a dataset comprising $\{100,200,400,800\}$ randomly selected documents for each language, extracted from the Wikipedia corpus. [^4] https://dumps.wikimedia.org/ Figure 4 shows the results of enhancement on high-resource languages (De, Fr, Zh, Es, Ru). The numbers represent the sizes of the training corpus when fine-tuning language-specific neurons, while "Random" represents the fine-tuning of an equivalent number of randomly chosen neurons using a corpus of $400$. Our findings reveal that fine-tuning with a few hundred documents yields significant performance improvements on multilingual tasks: $3.4\%$ on MGSM, $4.4\%$ on XQuAD, $4.3\%$ on X-CSQA, and $2.3\%$ on XLSum. Moreover, English performance is enhanced by an average of $3.7\%$ across all tasks. These results further confirm the effectiveness of MWork in interpreting structure functionality for LLM’s multilingual query handling, offering precise and independent methods for multilingual enhancement. When fine-tuning with 800 documents, the performance deteriorates compared to using 400 documents. This drop can be attributed to the incorporation of additional knowledge, which disrupts the original knowledge distribution and leads to overfitting of the model to Wikipedia. This can be addressed by mixing data from more sources such as textbooks or websites.
In addition, we verify the effectiveness of such enhancement method on low-resource languages, given that low-resource performance is relatively low with the original model. We select four languages including Vietnamese (Vi), Thai (Th), Arabic (Ar), and Swahili (Sw), covering languages with both latin and non-latin scripts and having corresponding testing set in our considered benchmarks. The model was then evaluated on four benchmarks, and the result shown in Table 4 is the average among tasks. It is evident that the fine-tuning method using language-specific neurons enhances the model’s multilingual performance in low-resource languages by an average of $2.2\%$. Notably, the improvement of $3.5\%$ in English performance is observed even without an English training corpus, indicating the effectiveness of the distinct language responsibilities assigned to neurons.
Figure 4: Enhancement results on high-resource languages, while the number is average among languages.
Table 6: Enhancement is achieved by fine-tuning Mistral-7b-v0.1 model utilizing $400$ documents from each language correspondingly. The results are averaged across four tasks. Performance on English (“En”) is obtained by averaging the results from three fine-tuned models.
Method En Vi Th Ar Sw
Original $41.1$ $32.7$ $25.6$ $21.7$ $15.1$
Random $40.8$ $32.7$ $25.2$ $21.2$ $15.1$
Lang-Spec $\mathbf{44.6}$ $\mathbf{34.9}$ $\mathbf{28.5}$ $\mathbf{23.4}$ $\mathbf{16.9}$
@@ 5 Related Work
In the era of Language and Linguistic Models (LLMs), numerous studies have been conducted to develop multilingual benchmarks (Zhang et al., 2023a), enhance multilingual performance without parameter adjustments through translation (Liang et al., 2023; Huang et al., 2023), aligning representations (Nguyen et al., 2023a; Salesky et al., 2023), prompting (Li et al., 2023b; Tanwar et al., 2023). Furthermore, certain works focus on improving multilingual abilities for a single task via cross-lingual transfer (Kim et al., 2017; Lin et al., 2019; Pfeiffer et al., 2020; Zhao et al., 2024b), while others aim to enhance multilingual proficiency by continuous training in one language to obtain mono-lingual LLMs (Cui et al., 2023a), or in multiple domain languages to obtain domain-lingual LLMs (Nguyen et al., 2023b). Additionally, some works achieve multilingual LLMs by training from scratch (Muennighoff et al., 2023). However, these studies are limited to specific task types or require substantial training corpora due to a lack of comprehensive understanding of the multilingual mechanisms of LLMs.
Conventional interpretability research investigates the significance of input features with their corresponding outputs (Vig, 2019; Hewitt and Liang, 2019; Qiu et al., 2020). In the era of LLMs, one brunch of work includes efforts to understand knowledge storage, with (Geva et al., 2021) initiating the study of the feed-forward layer as a knowledge base. Subsequent work has furthered this by altering neuron values (Dai et al., 2022b), mapping embeddings to words (Geva et al., 2022), modifying inputs to recover embeddings (Meng et al., 2022), and analyzing attention heads (Li et al., 2023a). Another line of research centers on the self-attention layer, examining its connection to reasoning capability (Hou et al., 2023; Stolfo et al., 2023; Friedman et al., 2023) by contrasting the reasoning tree based on attention weights.
@@ 6 Conclusion
In this work, we examine how LLMs handle multilingualism. The proposed multilingual workflow (MWork) suggests that LLMs initially understand queries by converting multilingual inputs into English, think in English in intermediate layers while incorporating multilingual knowledge, and generate responses aligned with the original language in the final layers. The validity of MWork is verified using Parallel Language-specific Neuron Detection (PLND), which identifies activated neurons for different languages without labeled data. By detecting language-specific neurons and fine-tuning them with a small training corpus, MWork enhances multilingual abilities in specific languages without compromising others, resulting in significant average improvements across tasks.
Limitation and Impact Statements
Our experiments are mainly conducted on models with a size of approximately 7 billion parameters (i.e., vicuna-7b-v1.5 and mistral-7b-v1.0), primarily due to computational resource constraints. Testing our methods on larger models could potentially yield additional insights, particularly in understanding the scalability of our proposed framework. Furthermore, our exploration into enhancing the multilingual abilities of language models through our framework was preliminary. Expanding our experiments to include a broader array of languages, especially those considered low-resource, could better demonstrate the effectiveness of our proposed framework. Moreover, extending the scope of our experiments to evaluate other capabilities such as reasoning and multilingual knowledge extraction with specific datasets could provide a more comprehensive picture of the potential benefits of our approach.
Our paper has the potential to significantly enhance the multilingual ability of LLMs. By effectively improving their performance across all languages, it can enable the development of real multilingual models that excel in various applications, promoting better communication and understanding among diverse linguistic communities.
@@ Appendix A English and Non-English Tokens
Appendix B Multilingual Corpus
Note that our selection criterion for the number of documents is based on achieving substantial coverage of each language’s vocabulary, ensuring that the selected contexts provide a representative sample of the language, as shown in Table 7.
Table 7: Corpus details across languages are tailored to encompass the majority of each language’s vocabulary, where “corpus size” indicates the number of contexts selected, “corpus vocab” represents the vocabulary coverage within the selected contexts, “vocab size” refers to the number of vocabularies of that language.
Language En De Fr Zh Es Ru
Corpus Size $180$k $30$k $50$k $20$k $20$k $20$k
Corpus Vocab $249$k $154$k $134$k $198$k $90$k $144$k
Vocab Size $273$k $148$k $135$k $329$k $93$k $150$k
Appendix C Interrelation of Language-Specific Neurons Across Languages
Using neurons identified by PLND, we investigate the relationships between languages via the degree of overlap among their language-specific neurons, defined as
(a) Mistral-7B-Instruct-v0.2.
(b) Vicuna-7b-v1.5.
We can observe that in both Mistral and Vicuna, the intersection with English from other languages is relatively limited, suggesting that English possesses a predominant number of language-specific neurons. Additionally, there is a pronounced tendency for languages belonging to the same family to demonstrate a higher degree of overlap with each other, such as Spanish, French, and English. Moreover, the feed-forward structure typically exhibits a higher degree of consistency in overlap across various languages, due to the shared world knowledge embedded within the neurons that is accessible to multiple languages.
Appendix D Prompts
Table 8 shows the zero-shot prompts for each dataset. Note that when conducting tests in other languages, prompts are translated into the respective languages.
Table 8: Zero-shot prompts for each dataset.
Task Zero-Shot Prompt
MGSM Let’s think step by step. Question: {question}
XQuAD {context} Question: {question}
XLSum Summarize the context in one sentence. Title: {title} Context: {article}
X-CSQA Question: {question}
Appendix E Original Performance
Table 9 shows the original performance of Vanilla and Mistral on four tasks.
Table 9: Assessing the baseline performance of Vicuna and Mistral across four representative multilingual tasks in selected languages, where Avg. is calculated among non-English languages.
Model Task En De Fr Zh Es Ru Avg.
Vicuna XQuAD $57.5$ $50.3$- $55.7$ $55.7$- $53.9$
MGSM $20.4$ $14.8$ $14.8$ $12.8$ $13.2$ $10.0$ $13.1$
X-CSQA $57.8$ $43.8$ $40.1$ $43.2$ $44.3$ $26.0$ $39.5$
XLSum $13.1$- $14.2$ $61.1$ $10.4$ $20.8$ $26.6$
Mistral XQuAD $57.1$ $48.5$- $64.3$ $54.1$- $55.6$
MGSM $46.0$ $21.2$ $26.0$ $31.6$ $31.2$ $21.6$ $26.3$
X-CSQA $61.7$ $40.0$ $40.4$ $47.1$ $45.7$ $14.1$ $37.5$
XLSum $13.5$- $15.2$ $56.4$ $10.6$ $21.0$ $25.8$
Appendix F Hyper-parameters